### v2

@@ A.4 $D\bar{B}^{*}\to D^{*}\bar{B}^{*}$
Figure 15: LO contact, NLO contact, LO OPE, NLO OPE and NLO TPE diagrams of the process $D\bar{B}^{*}\to D^{*}\bar{B}^{*}$.
@@
The one-loop corrections to the contact amplitudes are listed as:
@@
×$\times[\epsilon^{*\mu}_{3}(\epsilon_{2}\times\epsilon^{*}_{4})_{\mu}],$ (210)
@@
×$\times[\epsilon^{*\mu}_{3}(\epsilon_{2}\times\epsilon^{*}_{4})_{\mu}].$ (212)
@@
×$\times[\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu}],$ (219)
@@
×$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu},$ (220)
@@
×$\times[\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu}],$ (221)
@@
×$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu},$ (222)
@@
×$\times[\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu}],$ (223)
@@
×$\times[\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu}].$ (224)
In Tables 19-20, we list the coefficients of the above contact amplitudes at NLO and OPE amplitudes at NLO, respectively. The TPE amplitudes are given by:
@@
×$\times q_{\mu}q_{\nu}\epsilon^{*\mu}_{3}(\epsilon_{2}\times\epsilon^{*}_{4})^{\nu}\},$ (226)
@@
×$\times q_{\mu}q_{\nu}\epsilon^{*\mu}_{3}(\epsilon^{*}_{4}\times\epsilon_{2})^{\nu}\},$ (228)
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{a7.1}$ 00 00 00 00
$A_{a7.3(4)}$ 00 00 $0(\delta_{2})$ $\delta_{2}(0)$
$A_{a7.5(6)}$ 00 00 $-\delta_{1}$ $0(\delta_{1})$
$A_{a7.7}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta_{2}$ $-\delta_{1}$
$A_{a7.8}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $\delta_{2}$
$A_{a7.9}$ $\frac{1}{4}(D_{a}+E_{a}$ $+D_{b}+E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a}$ $+D_{b}-3E_{b})$ 00 $-\delta_{1}$
$A_{a7.10}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 00
$A_{a7.12}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ 00 $\delta_{1}$
$A_{a7.(13+14)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{1}$
$A_{a7.15}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $-\delta_{1}$ 00
$A_{a7.(16+17)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{2}$
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{b7.1(2)}$ $-\frac{1}{16}$ $\frac{3}{16}$ $0(\delta_{2})$ 00
$A_{b7.3(4)}$ $-\frac{1}{16}$ $\frac{3}{16}$ $-\delta_{1}$ $0(\delta_{1})$
$A_{b7.5(6)}$ $-\frac{1}{12}$ $\frac{1}{4}$ 00 00
$A_{b7.7}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b7.(8+9)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 $\delta_{1}$
$A_{b7.10}$ $\frac{1}{4}$ $-\frac{3}{4}$ $-\delta_{1}$ 00
$A_{b7.(11+12)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 $\delta_{2}$
$A_{b7.13(14)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b7.15(16)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{c7.2}$ $\frac{1}{16}$ $\frac{5}{16}$ $-\delta_{1}$ 00
$A_{c7.3}$ $\frac{1}{16}$ $\frac{5}{16}$ $-\delta_{1}$ $\delta_{2}$
$A_{c7.4}$ $\frac{9}{16}$ $-\frac{3}{16}$ $-\delta_{1}$ 00
$A_{c7.5}$ $\frac{9}{16}$ $-\frac{3}{16}$ $-\delta_{1}$ $\delta_{2}$
@@ A.5 $D^{*}\bar{B}\to D^{*}\bar{B}^{*}$
Figure 16: LO contact, NLO contact, LO OPE, NLO OPE and NLO TPE diagrams of the process $D^{*}\bar{B}\to D^{*}\bar{B}^{*}$.
@@
The one-loop corrections to the contact amplitude are listed as follows:
@@
×$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})_{\mu}],$ (246)
@@
×$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})_{\mu}].$ (248)
The one-loop corrections to the OPE amplitude are given by:
@@
×$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu}],$ (255)
@@
×$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu},$ (256)
@@
×$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu},$ (257)
@@
×$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu}],$ (258)
@@
×$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu}],$ (259)
@@
×$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu}].$ (260)
The coefficients appearing in the NLO contact and OPE amplitudes are listed in Table 22 and Table 23, respectively. And the TPE amplitudes:
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{a8.1}$ 00 00 $-\delta_{2}$ 00
$A_{a8.2(3)}$ 00 00 $-\delta_{2}(0)$ $\delta_{2}(0)$
$A_{a8.5(6)}$ 00 00 $0(\delta_{1})$ $\delta_{1}$
$A_{a8.7}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta$ $-\delta_{2}$
$A_{a8.8}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 $\delta_{1}$
$A_{a8.9}$ $\frac{1}{4}(D_{a}+E_{a}$ $+D_{b}+E_{b})$ $\frac{-3}{4}(D_{a}-3E_{a}$ $+D_{b}-3E_{b})$ 00 $-\delta_{2}$
$A_{a8.10}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 00 00
$A_{a8.12}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta_{2}$ 00
$A_{a8.(13+14)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{1}$
$A_{a8.(15+16)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 00 $\delta_{2}$
$A_{a8.(16+17)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $-\delta_{2}$ 00
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{b8.1(2)}$ $-\frac{1}{16}$ $\frac{3}{16}$ $-\delta_{2}$ $0(\delta_{2})$
$A_{b8.3(4)}$ $-\frac{1}{16}$ $\frac{3}{16}$ $0(\delta_{1})$ 00
$A_{b8.5(6)}$ $-\frac{1}{12}$ $\frac{1}{4}$ 00 00
$A_{b8.7}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b8.(8+9)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 $\delta_{1}$
$A_{b8.(10+11)}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 $\delta_{2}$
$A_{b8.12}$ $\frac{1}{4}$ $-\frac{3}{4}$ $-\delta_{2}$ 00
$A_{b8.13}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b8.14}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b8.15}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
$A_{b8.16}$ $\frac{1}{4}$ $-\frac{3}{4}$ 00 00
@@
$A$ $A$ $\omega_{1}$ $\omega_{2}$
$A_{c8.2}$ $\frac{1}{16}$ $\frac{5}{16}$ $-\delta_{2}$ 00
$A_{c8.3}$ $\frac{1}{16}$ $\frac{5}{16}$ $-\delta_{2}$ $\delta_{1}$
$A_{c8.4}$ $\frac{9}{16}$ $-\frac{3}{16}$ $-\delta_{2}$ 00
$A_{c8.5}$ $\frac{9}{16}$ $-\frac{3}{16}$ $-\delta_{2}$ $\delta_{1}$

### v3

@@ A.4 $D\bar{B}^{*}\to D^{*}\bar{B}^{*}$
Figure 15: LO contact ($a7$), LO OPE ($b7$), and NLO contact (a7.1a7.17a7.1-a7.17), NLO OPE (b7.1b7.16b7.1-b7.16), and NLO TPE (c7.1c7.5c7.1-c7.5) diagrams of the process $D\bar{B}^{*}\to D^{*}\bar{B}^{*}$.
@@
The one-loop corrections to the contact amplitudes are listed as
@@
$\times[\epsilon^{*\mu}_{3}(\epsilon_{2}\times\epsilon^{*}_{4})_{\mu}],$ (210)
@@
$\times[\epsilon^{*\mu}_{3}(\epsilon_{2}\times\epsilon^{*}_{4})_{\mu}].$ (212)
@@
$\times[\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu}],$ (219)
@@
$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu},$ (220)
@@
$\times[\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu}],$ (221)
@@
$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu},$ (222)
@@
$\times[\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu}],$ (223)
@@
$\times[\epsilon^{\mu}_{2}(\epsilon^{*}_{4}\times\epsilon^{*}_{3})^{\nu}].$ (224)
In Tables 19 and 20, we list the coefficients of the above contact amplitudes at NLO and OPE amplitudes at NLO, respectively. The TPE amplitudes are given by
@@
$\times q_{\mu}q_{\nu}\epsilon^{*\mu}_{3}(\epsilon_{2}\times\epsilon^{*}_{4})^{\nu}\},$ (226)
@@
$\times q_{\mu}q_{\nu}\epsilon^{*\mu}_{3}(\epsilon^{*}_{4}\times\epsilon_{2})^{\nu}\},$ (228)
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{a7.1}$ 0 0 0 0
$A_{a7.3(4)}$ 0 0 $0(\delta_{2})$ $\delta_{2}(0)$
$A_{a7.5(6)}$ 0 0 $\delta_{1}$ $0(\delta_{1})$
$A_{a7.7}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta_{2}$ $\delta_{1}$
$A_{a7.8}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 $\delta_{2}$
$A_{a7.9}$
$\frac{1}{4}(D_{a}+E_{a}$
$D_{b}+E_{b})$
$\frac{-3}{4}(D_{a}-3E_{a}$
$D_{b}-3E_{b})$
0 $\delta_{1}$
$A_{a7.10}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 0
$A_{a7.12}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ 0 $\delta_{1}$
$A_{a7.(13+14)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{1}$
$A_{a7.15}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $\delta_{1}$ 0
$A_{a7.(16+17)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{2}$
@@ 0 $\delta_{1}$
AA AA $\omega_{1}$ $\omega_{2}$
$A_{b7.1(2)}$ $\frac{1}{16}$ $\frac{3}{16}$ $0(\delta_{2})$ 0
$A_{b7.3(4)}$ $\frac{1}{16}$ $\frac{3}{16}$ $\delta_{1}$ $0(\delta_{1})$
$A_{b7.5(6)}$ $\frac{1}{12}$ $\frac{1}{4}$ 0 0
$A_{b7.7}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b7.(8+9)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 $\delta_{1}$
$A_{b7.10}$ $\frac{1}{4}$ $\frac{3}{4}$ $\delta_{1}$ 0
$A_{b7.(11+12)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 $\delta_{2}$
$A_{b7.13(14)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b7.15(16)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{c7.2}$ $\frac{1}{16}$ $\frac{5}{16}$ $\delta_{1}$ 0
$A_{c7.3}$ $\frac{1}{16}$ $\frac{5}{16}$ $\delta_{1}$ $\delta_{2}$
$A_{c7.4}$ $\frac{9}{16}$ $\frac{3}{16}$ $\delta_{1}$ 0
$A_{c7.5}$ $\frac{9}{16}$ $\frac{3}{16}$ $\delta_{1}$ $\delta_{2}$
@@ A.5 $D^{*}\bar{B}\to D^{*}\bar{B}^{*}$
Figure 16: LO contact ($a8$), LO OPE ($b8$), and NLO contact (a8.1a8.17a8.1-a8.17), NLO OPE (b8.1b8.16b8.1-b8.16), and NLO TPE (c8.1c8.5c8.1-c8.5) diagrams of the process $D^{*}\bar{B}\to D^{*}\bar{B}^{*}$.
@@
$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})_{\mu}],$ (246)
@@
$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})_{\mu}].$ (248)
The one-loop corrections to the OPE amplitudes are given by
@@
$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu}],$ (255)
@@
$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu},$ (256)
@@
$\times\frac{q_{\mu}q_{\nu}}{q^{2}-m^{2}}\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu},$ (257)
@@
$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu}],$ (258)
@@
$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu}],$ (259)
@@
$\times[\epsilon^{\mu}_{1}(\epsilon^{*}_{3}\times\epsilon^{*}_{4})^{\nu}].$ (260)
The coefficients appearing in the NLO contact and OPE amplitudes are listed in Tables 22 and 23, respectively, and the TPE amplitudes
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{a8.1}$ 0 0 $\delta_{2}$ 0
$A_{a8.2(3)}$ 0 0 $\delta_{2}(0)$ $\delta_{2}(0)$
$A_{a8.5(6)}$ 0 0 $0(\delta_{1})$ $\delta_{1}$
$A_{a8.7}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta$ $\delta_{2}$
$A_{a8.8}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 $\delta_{1}$
$A_{a8.9}$
$\frac{1}{4}(D_{a}+E_{a}$
$D_{b}+E_{b})$
$\frac{-3}{4}(D_{a}-3E_{a}$
$D_{b}-3E_{b})$
0 $\delta_{2}$
$A_{a8.10}$ $\frac{1}{4}(D_{b}+E_{b})$ $\frac{-3}{4}(D_{b}-3E_{b})$ 0 0
$A_{a8.12}$ $\frac{1}{4}(D_{a}+E_{a})$ $\frac{-3}{4}(D_{a}-3E_{a})$ $\delta_{2}$ 0
$A_{a8.(13+14)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{1}$
$A_{a8.(15+16)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ 0 $\delta_{2}$
$A_{a8.(16+17)}$ $D_{b}+E_{b}$ $D_{b}-3E_{b}$ $\delta_{2}$ 0
@@ 0 $\delta_{2}$
AA AA $\omega_{1}$ $\omega_{2}$
$A_{b8.1(2)}$ $\frac{1}{16}$ $\frac{3}{16}$ $\delta_{2}$ $0(\delta_{2})$
$A_{b8.3(4)}$ $\frac{1}{16}$ $\frac{3}{16}$ $0(\delta_{1})$ 0
$A_{b8.5(6)}$ $\frac{1}{12}$ $\frac{1}{4}$ 0 0
$A_{b8.7}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b8.(8+9)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 $\delta_{1}$
$A_{b8.(10+11)}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 $\delta_{2}$
$A_{b8.12}$ $\frac{1}{4}$ $\frac{3}{4}$ $\delta_{2}$ 0
$A_{b8.13}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b8.14}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b8.15}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
$A_{b8.16}$ $\frac{1}{4}$ $\frac{3}{4}$ 0 0
@@
AA AA $\omega_{1}$ $\omega_{2}$
$A_{c8.2}$ $\frac{1}{16}$ $\frac{5}{16}$ $\delta_{2}$ 0
$A_{c8.3}$ $\frac{1}{16}$ $\frac{5}{16}$ $\delta_{2}$ $\delta_{1}$
$A_{c8.4}$ $\frac{9}{16}$ $\frac{3}{16}$ $\delta_{2}$ 0
$A_{c8.5}$ $\frac{9}{16}$ $\frac{3}{16}$ $\delta_{2}$ $\delta_{1}$